OpenStudy (anonymous):

so, how many can solve sqrt{x} + y = 11 x + sqrt{y} = 7 sole this by process and steps..

OpenStudy (anonymous):

Just looking at this gives a logical answer of x=4, y=9

OpenStudy (ybarrap):

Without solving any equations, we can find this solution. x and y have to be square integers and x has to be less than 11 and y less than 7. The only square integers less than 11 is 9, 4, 1 and 0 and the only squares less than 7 are 4,1 and 0. In the 1st equation, if x was 9 then y should be 8, which isn't a perfect square. If x was 4 then y is 9, which IS a perfect square. To check, look at the 2nd equation, 4 + 3 is 7, so that must be the solution.

OpenStudy (anonymous):

any1 can guess d answr... I askd 2 solv it by step..

OpenStudy (ybarrap):

$$ \sqrt{x} + y = 11\implies\sqrt{x}=11-y\implies x=(11-y)^2=121-22y+y^2 \\ x + \sqrt{y} = 7\implies\sqrt{y}=7-x\implies y=(7-x)^2= 49-14x+x^2\\ $$ Add \(x+y\): $$ x+y=121-22y+y^2 + 49-14x+x^2\\ 0=121-22y-y+y^2 + 49-14x-x+x^2\\ 0=170-15x-23y+x^2+y^2\\ $$ Which is a circle of radius, \(\sqrt{\cfrac{37}{2}}\): $$ (x-\cfrac{15}{2})^2+(y-\cfrac{23}{2})^2=\cfrac{37}{2} $$ http://en.wikipedia.org/wiki/Circle#Equations Some integer solutions include: 4,9 4,14 5,8 5,15 10,8 10,15 11,9 11,14 Hope this helps.

OpenStudy (anonymous):

@ybarrap: Your answer is wrong. For example pair x=4, y=14 doesn't solve the original equations. Only pair x=4, y=9 solves the equations.